OpenStudy (anonymous):

9 + 4x = 3 Solve for x.

OpenStudy (moongazer):

-3/2

OpenStudy (anonymous):

x=-3/2

OpenStudy (anonymous):

9 + 4x = 3 4x = 3-9 4x = -6 x = -6/4 and simplfied to be -3/2

OpenStudy (anonymous):

If \(9+4x=3\), then we can subtract 9 from both sides, giving us \(9-9+4x=3-9\), which simplifies to \(4x=-6\). We can then divide each side by 4, giving us \(x={-6/4}\), which simplifies to \(x=-3/2\).